OpenStudy (anonymous):

what is the derivative of (e^2x)/[(x^2)+1]?

OpenStudy (anonymous):

just apply quotient rule and chain rule

OpenStudy (anonymous):

Take u=e^2x and v=(x^2 +1) and apply u/v rule for differentiation

OpenStudy (anonymous):

using the product rule and chain rule,it is {2(e^2x)/([x^2] +1)} - {2x(e^2x)/[([x^2]+1)]^2} u can simplify a little more